OpenStudy (anonymous):

Can anyone help me by explaining how this problem is solved? 6w^2 - 40w + 66

OpenStudy (anonymous):

Take 2 common: \[2(3w^2 - 20w + 33) \implies 2(3w^3 - 9w - 11w + 33) \implies 2(3w - 11 )(w-3)\]

OpenStudy (anonymous):

How did you get - 9w - 11w ?

OpenStudy (anonymous):

Splitting the middle term: -20w can be written as : -9w - 11w -9 and -11 will add up to give -20 again..

OpenStudy (anonymous):

But why did you pick 9 and 11?

OpenStudy (anonymous):

What did you get when you multiply 3 with 33???

OpenStudy (anonymous):

99?

OpenStudy (anonymous):

YEs.. We have to choose now two numbers that sum = -20 and product = +99 See, (-9)*(-11) = 99 (First Condition) -9 - 11 = -20 (second condition) That is why I chose that..

OpenStudy (anonymous):

Oh.. ok.

OpenStudy (anonymous):

How do I apply these steps when I do other similar problems?

OpenStudy (anonymous):

In this general equation: \[ax^2 + bx + c\] Firstly multiply a and c = a*c Now you have to chose two numbers that sum to give \(b\) and product to give \(ac\)..